OpenStudy (anonymous):

What is the answer to -6a+3=-3(2a-1) A:a=5 B:a=10 C: no solution D:idenity

OpenStudy (shamil98):

identity. -6a + 3 = -3(2a-1) -6a + 3 = -6a + 3 3=3 all real numbers work.

OpenStudy (anonymous):

Thank you